OpenStudy (anonymous):

Bob is driving along a straight and level road straight toward a mountain. At some point on his trip he measures the angle of elevation to the top of the mountain and finds it to be 25°11'. Find the height of the mountain to the nearest foot if Bob is 19,427.5 feet from the center of the mountain at the base. 9135 ft 9235 ft 91,351 ft 913,515 ft

OpenStudy (anonymous):



OpenStudy (anonymous):

tangent (25º11')=h/19427.5

OpenStudy (anonymous):

how do you finish solving ? @CliffSedge

OpenStudy (anonymous):

sorry im just confused

OpenStudy (anonymous):

Solve for h by multiplying both sides by 19427.5.

OpenStudy (anonymous):

so 19427.5*(25º11')=h/19427.5*19427.5 ???

OpenStudy (anonymous):

19427.5*tan(25º11')=(h/19427.5)*19427.5 Don't forget the tangent function.

OpenStudy (anonymous):

the answer is A @CliffSedge thank you

OpenStudy (anonymous):

Yes, 9135 feet is what I got too.